# Understanding Generalizability of Diffusion Models Requires Rethinking the Hidden Gaussian Structure

### v2

@@ Abstract
In this work, we study the generalizability of diffusion models by looking into the hidden properties of the learned score functions, which are essentially a series of deep denoisers trained on various noise levels. We observe that as diffusion models transition from memorization to generalization, their corresponding nonlinear diffusion denoisers exhibit increasing linearity. This discovery leads us to investigate the linear counterparts of the nonlinear diffusion models, which are a series of linear models trained to match the function mappings of the nonlinear diffusion denoisers. Surprisingly, these linear denoisers are approximately the optimal denoisers for a multivariate Gaussian distribution characterized by the empirical mean and covariance of the training dataset. This finding implies that diffusion models have the inductive bias towards capturing and utilizing the Gaussian structure (covariance information) of the training dataset for data generation. We empirically demonstrate that this inductive bias is a unique property of diffusion models in the generalization regime, which becomes increasingly evident when the model’s capacity is relatively small compared to the training dataset size. In the case that the model is highly overparameterized, this inductive bias emerges during the initial training phases before the model fully memorizes its training data. Our study provides crucial insights into understanding the notable strong generalization phenomenon recently observed in real-world diffusion models.
@@ 2 Preliminary
$\nabla_{\bm{x}}\log p(\bm{x};\sigma(t))=(\mathcal{D}_{\bm{\theta}}(\bm{x};\sigma(t))-\bm{x}(t))/\sigma^{2},$ (2)
@@
    which is essentially a softmax-weighted combination of the finite data points. As proved in [24, 17], such diffusion denoisers $\mathcal{D}_{\mathrm{M}}(\bm{x};\sigma(t))$ can only generate exact replicas of the training samples, therefore they have no generalizability.
@@
    where $\bm{\Sigma}=\bm{U}\bm{\Lambda}\bm{U}^{T}$ is the SVD of the empirical covariance matrix, with singular values $\bm{\Lambda}=\operatorname{diag}\left(\lambda_{1},\cdots,\lambda_{d}\right)$ and $\tilde{\bm{\Lambda}}_{\sigma(t)}=\operatorname{diag}\left(\frac{\lambda_{1}}{\lambda_{1}+\sigma(t)^{2}}\cdots\frac{\lambda_{d}}{\lambda_{d}+\sigma(t)^{2}}\right)$. With this linear Gaussian denoiser, as proved in [25], the sampling trajectory of the probabilistic ODE (1) has close form:
@@
As the training dataset size increases, diffusion models transition from the memorization regime—where they can only replicate its training images—to the generalization regime, where the they produce high-quality, novel images [17]. While memorization can be interpreted as an overfitting of diffusion models to the training samples, the mechanisms underlying the generalization regime remain less well-understood. This study aims to explore and elucidate the inductive bias that enables effective generalization in diffusion models.
@@ 3.2 Inductive Bias towards Learning the Gaussian Structures
•   Similarity in principle components. As shown in Figure 4(right), for a wide noise range ($\sigma(t)\in[0.116,80.0]$), the leading singular vectors of the linear weights $\bm{W}{\sigma(t)}$ (denoted $\bm{U}_{\text{\emph{Linear}}}$) align well with $\bm{U}$, the singular vectors of the Gaussian weights. [^4] For $\sigma(t)\in[0.116,80.0]$, the less well recovered singular vectors have singular values close to 00, whereas those corresponding to high singular values are well recovered. This implies that $\bm{U}$, representing the principal components of the training data, is effectively captured by the diffusion models. In the low-noise regime ($\sigma(t)\in[0.002,0.116]$), however, $\mathcal{D}_{\bm{\theta}}$ approximates the identity mapping, leading to ambiguous singular vectors with minimal impact on image generation. Further analysis of $\mathcal{D}_{\bm{\theta}}$’s behavior in the low-noise regime is provided in Appendices D and F.1.
@@ 4.1 Gaussian Structures Emerge when Model Capacity is Relatively Small
•  Decreasing model capacity promotes the emergence of Gaussian structure at fixed dataset sizes. Next, we investigate the impact of model scale by training diffusion models with varying channel sizes $[4,8,16,32,64,128]$, corresponding to $[64\text{k},251\text{k},992\text{k},4\text{M},16\text{M},64\text{M}]$ parameters, on a fixed training dataset of 1094 images. Figure 7(left) shows that in the intermediate noise variance regime ($\sigma(t)\in[0.116,20]$), the discrepancy between $\mathcal{D}_{\bm{\theta}}$ and $\mathcal{D}_{\mathrm{G}}$ decreases with decreasing model scale, indicating that Gaussian structure emerges in low-capacity models. Figure 7(right) demonstrates that this trend corresponds to a transition from data memorization to the generation of images exhibiting Gaussian structure. Here we note that smaller models lead to larger discrepancy between $\mathcal{D}_{\bm{\theta}}$ and $\mathcal{D}_{\mathrm{G}}$ in the high-noise regime. This phenomenon arises because diffusion models employ a bell-shaped noise sampling distribution that prioritizes intermediate noise levels, resulting in insufficient training at high noise variances, especially when model capacity is limited. (see more details in Section F.2).
@@ 5 Connection between Strong Generalizability and Gaussian Structure
To validate this hypothesis, we examine two diffusion models with channel size 128, trained on non-overlapping datasets S1 and S2 (35,000 images each). Figure 9(a) shows that images generated by these models (bottom) closely match those from their corresponding Gaussian models (top), highlighting the Gaussian structure’s role in strong generalization.
Comparing Figure 9(a)(top) and (b)(top), we observe that $\mathcal{D}_{\mathrm{G}}$ generates nearly identical images whether the Gaussian structure is calculated on a small dataset (1,094 images) or a much larger one (70,000 images). This similarity emerges because datasets of the same class can exhibit similar Gaussian structure (empirical covariance) with relatively few samples—just hundreds for FFHQ. Given the Gaussian structure’s critical role in generalization, small datasets may already contain much of the information for generalization, contrasting previous assertions in [20] that strong generalization requires training on datasets of substantial size (more than $10^{5}$ images). However, smaller datasets increase memorization risk, as shown in Figure 9(b). To mitigate this, as discussed in Section 4, we can either reduce model capacity or implement early stopping (Figure 9(c)). Indeed, models trained on 1094 and 35000 images generate remarkably similar images, though the smaller dataset yields lower perceptual quality. This similarity further demonstrates that small datasets contain substantial generalization-relevant information closely tied to Gaussian structure. Further discussion on the connections and differences between our work and [20] are detailed in Appendix H.
@@ 6 Discussion
In this study, we empirically demonstrate that diffusion models in the generalization regime have the inductive bias towards learning diffusion denoisers that are close to the the corresponding linear Gaussian denoisers—the optimal solution under Gaussian data assumptions. While real-world image distributions are significantly different from Gaussian, our findings imply that diffusion models have the bias towards learning and utilizing low-dimensional data structures, such as the data covariance, for image generation. However, the underlying mechanism by which the nonlinear diffusion models, trained with gradient descent, exhibit such linearity remains unclear and warrants further investigation.
@@ H.2 GAHB Emerge only in Intermediate-Noise Regime
  • High-noise regime [10,80]. In this regime, the leading singular vectors [^6] We only care about leading singular vectors since the Jacobians in this regime are highly low-rank. The less well aligned singular vectors have singular values near 0. of the Jacobian matrix $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ well align with those of the Gaussian weights (the leading principle components of the training dataset), consistent with our finding that diffusion denoisers approximate linear Gaussian denoisers in this regime. Notice that DAEs trained sufficiently on separate noise levels (Figure 29) show stronger alignment compared to vanilla diffusion models (Figure 28), which suffer from insufficient training at high noise levels.
@@
Notice that the leading singular vectors of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ are the input directions that lead to the maximum variation in denoised outputs, thus revealing meaningful information on the local properties of $\mathcal{D}(\bm{x}_{t};\sigma(t))$ at $\bm{x}_{t}$. As demonstrated in Figure 30, perturbing input $\bm{x}_{t}$ along these vectors at difference noise regimes leads to distinct effects on the final generated images: (i) in the high-noise regime where the leading singular vectors align with the principle components of the training dataset, perturbing $\bm{x}_{t}$ along these directions leads to canonical changes such as image class, (ii) in the intermediate-noise regime where the GAHB emerge, perturbing $\bm{x}_{t}$ along the leading singular vectors modify image details such as colors while preserving overall image structure and (iii) in the low-noise regime where the leading singular vectors have no significant pattern, perturbing $\bm{x}_{t}$ along these directions yield no meaningful semantic changes.
@@
Figure 28: Evolution of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ across varying noise levels. Figure(a) shows the generation trajectory. Figure(b) shows the correlation matrix between Jacobian singular vectors $\bm{U}(\bm{x}_{t})$ and training dataset principal components $\bm{U}$. Notice that the leading singular vectors of $\bm{U}(\bm{x}_{t})$ and $\bm{U}$ well align in early timesteps but diverge in later timesteps. Figure(c) shows the first three principle components of the training dataset while figures(d-f) show the evolution of Jacobian’s first three singular vectors across noise levels. These singular vectors initially match the principle components but progressively adapt to input image geometry, before losing distinct patterns at very low noise levels. While we present only left singular vectors, right singular vectors exhibit nearly identical behavior and yield equivalent results.
Figure 29: Evolution of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ across varying noise levels for DAEs. We repeat the experiments in Figure 28 on DAEs that are sufficiently trained on each discrete noise levels. Notice that with sufficient training, the Jacobian singular vectors $\bm{U}(\bm{x}_{t})$ show a better alignment with principle components $\bm{U}$ in early timesteps.

### v5

@@ Abstract
In this work, we study the generalizability of diffusion models by looking into the hidden properties of the learned score functions, which are essentially a series of deep denoisers trained on various noise levels. We observe that as diffusion models transition from memorization to generalization, their corresponding nonlinear diffusion denoisers exhibit increasing linearity. This discovery leads us to investigate the linear counterparts of the nonlinear diffusion models, which are a series of linear models trained to match the function mappings of the nonlinear diffusion denoisers. Interestingly, these linear denoisers are nearly optimal for multivariate Gaussian distributions defined by the empirical mean and covariance of the training dataset, and they effectively approximate the behavior of nonlinear diffusion models. This finding implies that diffusion models have the inductive bias towards capturing and utilizing the Gaussian structure (covariance information) of the training dataset for data generation. We empirically demonstrate that this inductive bias is a unique property of diffusion models in the generalization regime, which becomes increasingly evident when the model’s capacity is relatively small compared to the training dataset size. In the case where the model is highly overparameterized, this inductive bias emerges during the initial training phases before the model fully memorizes its training data. Our study provides crucial insights into understanding the notable strong generalization phenomenon recently observed in real-world diffusion models.
@@ 2 Preliminary
$\nabla_{\bm{x}}\log p(\bm{x};\sigma(t))=(\mathcal{D}_{\bm{\theta}}(\bm{x};\sigma(t))-\bm{x})/\sigma(t)^{2},$ (2)
@@
    which is essentially a softmax-weighted combination of the finite data points. As proved in [24], such diffusion denoisers $\mathcal{D}_{\mathrm{M}}(\bm{x};\sigma(t))$ can only generate exact replicas of the training samples, therefore they have no generalizability.
@@
    where $\bm{\Sigma}=\bm{U}\bm{\Lambda}\bm{U}^{T}$ is the SVD of the empirical covariance matrix, with singular values $\bm{\Lambda}=\operatorname{diag}\left(\lambda_{1},\cdots,\lambda_{d}\right)$ and $\tilde{\bm{\Lambda}}_{\sigma(t)}=\operatorname{diag}\left(\frac{\lambda_{1}}{\lambda_{1}+\sigma(t)^{2}},\cdots,\frac{\lambda_{d}}{\lambda_{d}+\sigma(t)^{2}}\right)$. With this linear Gaussian denoiser, as proved in [25], the sampling trajectory of the probabilistic ODE (1) has close form:
@@
As the training dataset size increases, diffusion models transition from the memorization regime—where they can only replicate its training images—to the generalization regime, where the they produce high-quality, novel images [17]. While memorization can be interpreted as an overfitting of diffusion models to the training samples, the mechanisms underlying the generalization regime remain less well understood. This study aims to explore and elucidate the inductive bias that enables effective generalization in diffusion models.
@@ 3.2 Inductive Bias towards Learning the Gaussian Structures
•   Similarity in principal components. As shown in Figure 4(right), for a wide noise range ($\sigma(t)\in[0.116,80.0]$), the leading singular vectors of the linear weights $\bm{W}_{\sigma(t)}$ (denoted $\bm{U}_{\text{\emph{Linear}}}$) align well with $\bm{U}$, the singular vectors of the Gaussian weights. [^4] For $\sigma(t)\in[0.116,80.0]$, the less well recovered singular vectors have singular values close to 00, whereas those corresponding to high singular values are well recovered. This implies that $\bm{U}$, representing the principal components of the training data, is effectively captured by the diffusion models. In the low-noise regime ($\sigma(t)\in[0.002,0.116]$), however, $\mathcal{D}_{\bm{\theta}}$ approximates the identity mapping, leading to ambiguous singular vectors with minimal impact on image generation. Further analysis of $\mathcal{D}_{\bm{\theta}}$’s behavior in the low-noise regime is provided in Appendices D and F.1.
@@ 4.1 Gaussian Structures Emerge when Model Capacity is Relatively Small
•  Decreasing model capacity promotes the emergence of Gaussian structure at fixed dataset sizes. Next, we investigate the impact of model scale by training diffusion models with varying channel sizes $[4,8,16,32,64,128]$, corresponding to $[64\text{k},251\text{k},992\text{k},4\text{M},16\text{M},64\text{M}]$ parameters, on a fixed training dataset of 1094 images. Figure 7(left) shows that in the intermediate noise variance regime ($\sigma(t)\in[0.116,20]$), the discrepancy between $\mathcal{D}_{\bm{\theta}}$ and $\mathcal{D}_{\mathrm{G}}$ decreases with decreasing model scale, indicating that Gaussian structure emerges in low-capacity models. Figure 7(right) demonstrates that this trend corresponds to a transition from data memorization to the generation of images exhibiting Gaussian structure. Here we note that smaller models lead to larger discrepancy between $\mathcal{D}_{\bm{\theta}}$ and $\mathcal{D}_{\mathrm{G}}$ in the high-noise regime. This phenomenon arises because diffusion models employ a bell-shaped noise sampling distribution that prioritizes intermediate noise levels, resulting in insufficient training at high noise variances, especially when model capacity is limited (see more details in Section F.2).
@@ 5 Connection between Strong Generalizability and Gaussian Structure
To validate this hypothesis, we examine two diffusion models with channel size 128, trained on non-overlapping datasets S1 and S2 (35000 images each). Figure 9(a) shows that images generated by these models (bottom) closely match those from their corresponding Gaussian models (top), highlighting the Gaussian structure’s role in strong generalization.
Comparing Figure 9(a)(top) and (b)(top), we observe that $\mathcal{D}_{\mathrm{G}}$ generates nearly identical images whether the Gaussian structure is calculated on a small dataset (1094 images) or a much larger one (35000 images). This similarity emerges because datasets of the same class can exhibit similar Gaussian structure (empirical covariance) with relatively few samples—just hundreds for FFHQ. Given the Gaussian structure’s critical role in generalization, small datasets may already contain much of the information needed for generalization, contrasting previous assertions in [20] that strong generalization requires training on datasets of substantial size (more than $10^{5}$ images). However, smaller datasets increase memorization risk, as shown in Figure 9(b). To mitigate this, as discussed in Section 4, we can either reduce model capacity or implement early stopping (Figure 9(c)). Indeed, models trained on 1094 and 35000 images generate remarkably similar images, though the smaller dataset yields lower perceptual quality. This similarity further demonstrates that small datasets contain substantial generalization-relevant information closely tied to Gaussian structure. Further discussion on the connections and differences between our work and [20] are detailed in Appendix H.
@@ 6 Discussion
In this study, we empirically demonstrate that diffusion models in the generalization regime have the inductive bias towards learning diffusion denoisers that are close to the corresponding linear Gaussian denoisers. Although real-world image distributions are significantly different from Gaussian, our findings imply that diffusion models have the bias towards learning and utilizing low-dimensional data structures, such as the data covariance, for image generation. However, the underlying mechanism by which the nonlinear diffusion models, trained with gradient descent, exhibit such linearity remains unclear and warrants further investigation.
@@ H.2 GAHB Emerge only in Intermediate-Noise Regime
  • High-noise regime [10,80]. In this regime, the leading singular vectors [^6] We only care about leading singular vectors since the Jacobians in this regime are highly low-rank. The less well aligned singular vectors have singular values near 0. of the Jacobian matrix $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ well align with those of the Gaussian weights (the leading principal components of the training dataset), consistent with our finding that diffusion denoisers approximate linear Gaussian denoisers in this regime. Notice that DAEs trained sufficiently on separate noise levels (Figure 29) show stronger alignment compared to vanilla diffusion models (Figure 28), which suffer from insufficient training at high noise levels.
@@
Notice that the leading singular vectors of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ are the input directions that lead to the maximum variation in denoised outputs, thus revealing meaningful information on the local properties of $\mathcal{D}(\bm{x}_{t};\sigma(t))$ at $\bm{x}_{t}$. As demonstrated in Figure 30, perturbing input $\bm{x}_{t}$ along these vectors at difference noise regimes leads to distinct effects on the final generated images: (i) in the high-noise regime where the leading singular vectors align with the principal components of the training dataset, perturbing $\bm{x}_{t}$ along these directions leads to canonical changes such as image class, (ii) in the intermediate-noise regime where the GAHB emerge, perturbing $\bm{x}_{t}$ along the leading singular vectors modify image details such as colors while preserving overall image structure and (iii) in the low-noise regime where the leading singular vectors have no significant pattern, perturbing $\bm{x}_{t}$ along these directions yield no meaningful semantic changes.
@@
Figure 28: Evolution of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ across varying noise levels. Figure(a) shows the generation trajectory. Figure(b) shows the correlation matrix between Jacobian singular vectors $\bm{U}(\bm{x}_{t})$ and training dataset principal components $\bm{U}$. Notice that the leading singular vectors of $\bm{U}(\bm{x}_{t})$ and $\bm{U}$ well align in early timesteps but diverge in later timesteps. Figure(c) shows the first three principal components of the training dataset while figures(d-f) show the evolution of Jacobian’s first three singular vectors across noise levels. These singular vectors initially match the principal components but progressively adapt to input image geometry, before losing distinct patterns at very low noise levels. While we present only left singular vectors, right singular vectors exhibit nearly identical behavior and yield equivalent results.
Figure 29: Evolution of $\nabla\mathcal{D}(\bm{x}_{t};\sigma(t))$ across varying noise levels for DAEs. We repeat the experiments in Figure 28 on DAEs that are sufficiently trained on each discrete noise levels. Notice that with sufficient training, the Jacobian singular vectors $\bm{U}(\bm{x}_{t})$ show a better alignment with principal components $\bm{U}$ in early timesteps.